OpenStudy (anonymous):

Let y vary inversely as the second power of x. If x doubles, what happens to y?

OpenStudy (anonymous):

y vary inversely as the second power of x y=k/x^2, where k is any constant k = x^2 y x doubles, so x becomes 2x k = (2x)^2 y = 4x^2y y=k/x^2 becomes y = 4x^2y / x^2 = 4y y quadruples.